OpenStudy (anonymous):

The speed of a stream is 3mph. a boat travels 11 miles upstream in the same time it takes to travel 17 miles downstream. what is the speed of the boat in still water?

OpenStudy (radar):

Let x = speed of boat in still water x-3 = speed up stream x+3 = speed down stream d=speed times time 11= (x-3) t 17=(x+3) t t=11/(x-3) t=17/(x+3) t=t so 11/(x-3)=17/(x+3) cross multiply getting 11(x+3) = 17(x-3) 11x +33= 17x-51 16x=18 x=18/16=9/8 mph